A consumer has a demand function for rice of the form:





X = 3 + m/2p





Assume that his/her income per month that is allocated to consumption of rice in a month is kshs 600/-The price of one kg of rice is kshs 100/-. During harvesting, the price of rice falls by shs 25 per kg.





Compute the consumption of rice by the consumer in a month before and after the fall in price.





Compute the Substitution Effect on the consumer





Compute the Income Effect





What is the total effect of the price change?





A consumer is faced with the Utility function of the form:





U(x,y) = αLogx + βLogy





If this Utility function is represented by a map of indifference curves, find the slope of a particular indifference curve at any point.

1
Expert's answer
2021-12-06T17:12:06-0500

(a)

Consumption before:

"x=3+\frac{m}{2p}"

"=3+\frac{600}{2\times 100}=6"

Consumption after:

"x=3+\frac{m}{2p}"

"=3+\frac{600}{2\times75}=7"

(b)

Substitution Effect:

"=x_h-x_0"

"=3+\frac{600\times (\frac{75}{100})}{2\times 75}=3"

(c)

Income Effect:

"=x_1-x_h"

"=3+\frac{600}{2\times75}-3+\frac{(600\times\frac{75}{100})}{2\times 75}=1"

(d)

Total Effect:

"=3+\frac{600}{(2\times100)}-3+\frac{600}{(2\times75)}=-1"